A body has a positive charge of 8*10^-19c.it has


a)An excess of 5 electron      b)a deficiency of 5 electrons


c)an excess of 8 electron       d)a deficiency of 8 electrons


 

Grade:

3 Answers

SOHAN SARANGI
12 years ago

The answer is that the body has deficiency of 5 electrons. Because the positve charge is formed when electrons are given away from the body. Never the positive charges are transferred..

Also 8*10^-19C is a multiple of 5*charge of 1 electron..

Hence (b) is the correct option.

nikhil arora
12 years ago

it shud be (b).

because there is a positive charge , the body shud have less electron than protons and thus there is deficiency. and there is 1.6 * 10-19 C charge on each electron so there are 5 less electrons.

SHANTANU SHUKLA
12 years ago

body has charge Q=8*10-19  SO FROM    Q=Ne     n=q/e      8*10-19/1.6*10-19   =5  deficiency electrons